1.- Sara is a businesswoman who has established a web page where people can design and buy a t-shirt. Sara pays 1000 per week for the web server and the internet connection. The shirts that his clients design are sent to another company and Sara pays this company 20 per shirt. Sara has no other costs. The following table shows your demand scheme for Sara's shirts:

Price (dollars per sweatshirt) Quantity demanded (sweatshirts per week) 100 20 80 80 100

a) Calculate what will be the price, production and economic benefits that maximize Sara's profits.
b) Do you expect other companies to enter the t-shirt business online and compete with Sara?
c) What happens to the demand for Sara's web shirts in the long term? What happens to Sara's economic profits in the long term?
Answer #1

(a)

TR = P x Q

MR = Change in TR / Change in Q

P Q TR MR
0 100 0
20 80 1600 -80
40 60 2400 -40
60 40 2400 0
80 20 1600 40
100 0 0 80

Profit is maximized when MR = MC = 20.

When Q = 40, MR < MC and when Q = 20, MR > MC.

So Q = 20 and P = 80.

Profit = Q x (P - MC) - TFC = 20 x (80 - 20) - 1,000 = (20 x 60) - 1,000 = 1,200 - 1,000 = 200

(b)

Since profit > 0, it will attract entry of new firms.

(c)

In long term, higher number of firms will decrease demand for Sara's firm and this will continue until Sara earns zero economic profit.